OpenStudy (anonymous):

Mandy and 2 friends bought pencils. They divided some equally amongst themselves and have her brother 3. There were 19 pencils left. Solve the equation p/3-3=19 to f find the total number of pencils they bought.

OpenStudy (anonymous):

p/3-3=19 p/3= 22 p= 7.3

OpenStudy (anonymous):

@Jessicaisolda196 im so sorry that your child had to see all the things that he was saying

OpenStudy (anonymous):

Thank you mondona. That kid is a moron.

OpenStudy (anonymous):

sike stil here

OpenStudy (mathmale):

I'd suggest you let n represent the total number of pencial purchased by the 3 students. We are to find the value of n (that is, we are to find the total number of pencils they bought. In the end they had 19 pencils left. Can you translate this and all the other givens into an equation?

OpenStudy (mathmale):

Solve the equation p/3-3=19. You might find this easier if you'd type it in (using Equation Editor) as \[\frac{ p }{ 3 }-3=19\]